OpenStudy (anonymous):

evaluate sum from n=1 to infinity ((3n-1)/(4n+3))^(2n) by the root test, is it divergent, convergent or inconclusive?

OpenStudy (anonymous):

converges for sure. that \(^{2n}\) in the denominator makes the terms go to zero lickety split

OpenStudy (anonymous):

\[\lim_{n\to \infty}\left(\frac{3n-1}{(4n+3)^{2n}}\right)^{\frac{1}{n}}\] \[\lim_{n\to \infty}\left(\frac{(3n+1)}{(4n+3)^2}\right)=0\]

OpenStudy (anonymous):

oops i made a mistake, should be a \(\frac{1}{n}\) in the numerator as well

OpenStudy (anonymous):

the answer was 9/16...

OpenStudy (anonymous):

\[ \left(\left(\frac{3 n-1}{4 n+3}\right)^{2 n}\right)^{\frac{1}{n}}=\frac{(3 n-1)^2}{(4 n+3)^2} \] For n near infinity the above ration behaves like \[ \frac { 9n^2}{16 n^2}= \frac 9 {16} \]

OpenStudy (anonymous):

To see it in more steps, \[ \frac{(3 n-1)^2}{(4 n+3)^2}= \frac {9 n^2-6 n+1 } {16 n^2+24 n+9} = \frac { \frac {9 n^2-6 n+1 } {n^2} } {\frac{16 n^2+24 n+9}{n^2}}=\frac { 9 - \frac 6 n + \frac 1 {n^2}} {16 + \frac {24}{n}+ \frac 9{n^2} } \] you can see that the ratio tends to 9/16 when n goes to infinity.

OpenStudy (anonymous):

Is it convergent or divergent?